# Fermionic Partial Transpose in the Overlap Matrix Framework for Entanglement Negativity

### v2

@@ Appendix B The Fisher-Hartwig conjecture
$\frac{\text{d}\ln\ D_{L}(\lambda)}{\text{d}\lambda}$ $=L\left[\frac{1}{\lambda-1}-\frac{k_{F}}{2\pi}\left(\frac{1}{\lambda-1}-\frac{1}{\lambda}\right)\right]$ (41)
$\hskip 10.00002pt-\left(2\ln\ L+(1+\gamma_{E})+\sum_{n=1}^{\infty}\left[\frac{n}{n^{2}-\beta^{2}(\lambda)}-\frac{1}{n}\right]+\ln\ (2-2\cos k_{F})\right)\frac{\text{d}\beta^{2}(\lambda)}{\text{d}\lambda}$
$=L\left[\frac{1}{\lambda-1}-\frac{k_{F}}{2\pi}\left(\frac{1}{\lambda-1}-\frac{1}{\lambda}\right)\right]$
$\hskip 10.00002pt-\frac{2\beta(\lambda)}{\pi\text{i}\lambda(\lambda-1)}\left[\ln\ L+(1+\gamma_{E})+\ln\ (2|\sin\frac{1}{2}k_{F}|)+\Upsilon(\lambda)\right],$
where $\Upsilon(\lambda)=\sum_{n=1}^{\infty}\frac{n}{n^{2}-\beta^{2}(\lambda)}-\frac{1}{n}$. For the linear term in $L$, since all of them exist a first-order pole in $\lambda$, the integral over $\lambda$ ultimately contributes zero to them based on the residue theorem. Therefore, the entanglement quantity is given by
Entanglement $\text{Entanglement quantity}=\frac{1}{\pi^{2}}\oint\text{d}\lambda f(\lambda)\frac{\beta(\lambda)}{\lambda(\lambda-1)}\left[\ln\ L+(1+\gamma_{E})+\ln\ (2|\sin\frac{1}{2}k_{F}|)+\Upsilon(\lambda)\right],$ (42)
where the integral contour is shown in Fig. S2. The residue theorem ensures that the contributions from the integration paths $\overrightarrow{FED}$ and $\overrightarrow{CAB}$ vanish. Consequently, the integral simplifies to
Entanglement quantity $=\frac{1}{\pi^{2}}\left(\int_{1+\text{i}0^{+}}^{0+\text{i}0^{+}}+\int_{0+\text{i}0^{-}}^{1+\text{i}0^{-}}\right)\text{d}\lambda f(\lambda)\frac{\beta(\lambda)}{\lambda(\lambda-1)}$ (43)
$\hskip 10.00002pt\left[\ln\ L+(1+\gamma_{E})+\ln\ (2|\sin\frac{1}{2}k_{F}|)+\Upsilon(\lambda)\right].$
In the complex number field the phase of $\beta\left(\lambda\right)$ must be taken into account to ensure a consistent and well-defined formulation
$\beta(x(\in\mathbb{R})+\mathrm{i}0^{\pm})$ $=\frac{1}{2\pi\mathrm{i}}\left(\mathrm{Ln}\frac{\left(x-1+\mathrm{i}0^{\pm}\right)\left(x-\mathrm{i}0^{\pm}\right)}{x^{2}}\right)$ (44)
@@
Figure S2: The integral contour in Eq. (42) consists of arcs $\overrightarrow{FED}$ and $\overrightarrow{CBA}$ with radii $\epsilon\to 0$. The midpoint of these arcs originate from the range of eigenvalues of the overlap matrix. The imaginary part of $\overrightarrow{AF}$($\overrightarrow{DC}$) is given by $\delta=0^{+}\left(-\delta=0^{-}\right)$.
@@
$\hskip 10.00002pt+\frac{1}{\pi^{2}}\sum_{n=1}^{\infty}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}n\left[\frac{\text{i}W(x)-\frac{1}{2}}{n^{2}-[-\text{i}W(x)+\frac{1}{2}]^{2}}\right.$
$\hskip 70.0001pt\left.+\frac{-\text{i}W(x)-\frac{1}{2}}{n^{2}-[-\frac{1}{2}-\text{i}W(x)]^{2}}+\frac{1}{n^{2}}\right]$
@@
$E_{2}$ $=\frac{1}{\pi^{2}}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}\left(-1\right)\left[-\gamma_{E}+\sum_{n=0}^{\infty}\frac{1}{n+1}-\sum_{n=1}^{\infty}\frac{1}{n}\right.$ (47)
$\hskip 130.0002pt\left.-\frac{1}{2}\left(\psi(\frac{1}{2}-\text{i}W(x))+\psi(\frac{1}{2}+\text{i}W(x))\right)-1\right]$
$=\frac{-1}{\pi^{2}}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}\left(-\gamma_{E}-1\right)$
@@ Appendix C The eigenvalues of the density matrix after partial transpose in the overlap matrix approach
In this appendix, we will generalize our result to a random hopping model, i.e.
$H=\sum_{i,j}t_{ij}c_{i}^{{\dagger}}c_{j}=\bm{c}^{{\dagger}}t\bm{c},$ (49)
where the elements of $t$ are random numbers following a uniform distribution while maintaining Hermiticity. This can be achieved by generating random matrices $T_{1}$ and $T_{2}$ to construct $T=T_{1}+\text{i}T_{2}$, and then symmetrizing as $t=\frac{T+T^{{\dagger}}}{2}$. Other types of random $t$ can also be generated to further validate our results. Numerical findings confirm that the overlap matrix approach aligns with exact diagonalization, as shown in Fig. S3.
Figure S3: Comparison of the eigenvalues of $\rho^{T_{B}^{f}}$ and $\tilde{\rho}^{T_{B}^{f}}$: (a) the congruence of eigenvalues between $\rho^{T_{B}^{f}}$ and $\tilde{\rho}^{T_{B}^{f}}$; (b) difference in real parts of eigenvalues; (c) difference in imaginary parts of eigenvalues. We select $N=12$ with an even bipartite geometry. The results indicate that the difference is on the order of $10^{-15}$, demonstrating near-perfect agreement.
Appendix D A proof of rationality of partial trace in the overlap matrix approach
In this appendix, we expand the Eq. (10) into Fock space and compare with the reduced density matrix partially traced in the original basis. We prove that these two equations are completely identical.
@@ Appendix D A proof of rationality of partial trace in the overlap matrix approach
|$\ket{\Psi}$ $=\prod_{\alpha=1}^{M}f_{\alpha}^{\dagger}|0\rangle$ (50)
$=\prod_{\alpha=1}^{M}\sum_{i=1}^{N}\left(c_{i}^{\dagger}U_{i,\alpha}\right)\ket{0}$
@@
$=\sum_{1\leq j_{1}<j_{2}<\cdots<j_{M}\leq N}\left(\sum_{\sigma\in S_{M}}\mathrm{sgn}(\sigma)U_{j_{\sigma(1)},1}U_{j_{\sigma(2)},2}\cdots U_{j_{\sigma(M)},M}\right)c_{j_{1}}^{\dagger}c_{j_{2}}^{\dagger}\cdots c_{j_{M}}^{\dagger}|0\rangle$
@@
where in the fourth line $\sum_{\sigma\in S_{M}}$ represents the summation over all elements of the $M$th-order symmetric group $S_{M}$ ($M$ denotes the paticle number) and in the last line $J$ implies a permitted sequence of $\left\{j_{1},j_{2},\cdots,j_{M}\right\}$. Accordingly the density matrix
@@
The partial trace operation ensures that the particles in region $B$ remain the same in both bra and ket states, i.e. $J_{B}=K_{B}$. Then we decompose the numerous terms in $\rho_{A}$ based on the particles in region $A$
$\rho_{A}=\sum_{m=0}^{\min(M,|A|)}\sum_{\begin{subarray}{c}J_{A},K_{A}\subseteq A\\ |J_{A}|=|K_{A}|=m\end{subarray}}\left(\sum_{\begin{subarray}{c}J_{B}\subseteq B\\ |J_{B}|=M-m\end{subarray}}\det(U_{J_{A}\cup J_{B}})\det(U_{K_{A}\cup J_{B}})^{*}\right)|J_{A}\rangle\langle K_{A}|.$ (52)
It should be noted that the maximum particle number in region $A$ is $\min\left(M,\left|A\right|\right)$, where $\left|A\right|$ represents the the number of sites in region $A$. From Sec. II we also know that rotating $\left\{f_{\alpha}\right\}_{\alpha=1}^{M}$ to $\left\{d_{Ai}\right\}_{i=1}^{M}$ and $\left\{d_{Bi}\right\}_{i=1}^{M}$ does not alter the density matrix. Hence the reduced density matrix can also be expressed as
$\rho_{A}=\sum_{m=0}^{\min(M,|A|)}\sum_{\begin{subarray}{c}J_{A},K_{A}\subseteq A\\ |J_{A}|=|K_{A}|=m\end{subarray}}\left(\sum_{\begin{subarray}{c}J_{B}\subseteq B\\ |J_{B}|=M-m\end{subarray}}\det((U\mathcal{U})_{J_{A}\cup J_{B}})\det((U\mathcal{U})_{K_{A}\cup J_{B}})^{*}\right)|J_{A}\rangle\langle K_{A}|.$ (53)
Then we expand the Eq. (10):
@@
$=\prod_{\gamma=1}^{M}\left(P_{\gamma}\frac{\sum_{\alpha=1}^{M}\sum_{j\in A}c_{j}^{\dagger}U_{j,\alpha}\mathcal{U}_{\alpha,\gamma}}{\sqrt{P_{\gamma}}}|0\rangle\langle 0|\frac{\sum_{\beta=1}^{M}\sum_{j\in A}c_{j}U_{j,\beta}^{*}\mathcal{U}_{\beta,\gamma}^{*}}{\sqrt{P_{\gamma}}}+(1-P_{\gamma})|0\rangle\langle 0|\right)$
@@
where the symbol $\tilde{\rho}_{A}$ is introduced to distinguish it from the original reduced density matrix. In the fourth line we express $1-P_{\gamma}$ as $\sum_{j\in B}\sum_{\alpha,\beta=1}^{M}U_{j,\alpha}\mathcal{U}_{\alpha,\gamma}U_{j,\beta}^{*}\mathcal{U}_{\beta,\gamma}^{*}$, which can be verified by the identity $\left(\mathcal{U}^{\dagger}M_{B}\mathcal{U}\right)_{\alpha\beta}=\delta_{\alpha\beta}\left(1-P_{\alpha}\right)$. With the essential identity
00 0 $=\sum_{\mu\nu}\sum_{i\in B}\mathcal{U}_{\mu\beta}^{*}U_{i\mu}^{*}U_{i\nu}\mathcal{U}_{\nu\alpha}$ (55)
@@
which help derive the equation below
$\tilde{\rho}_{A}=\sum_{m=0}^{\min(M,|A|)}\sum_{\begin{subarray}{c}J_{A},K_{A}\subseteq A\\ |J_{A}|=|K_{A}|=m\end{subarray}}\left(\sum_{\begin{subarray}{c}J_{B}\subseteq B\\ |J_{B}|=M-m\end{subarray}}\det((U\mathcal{U})_{J_{A}\cup J_{B}})\det((U\mathcal{U})_{K_{A}\cup J_{B}})^{*}\right)|J_{A}\rangle\langle K_{A}|,$ (56)
from the Eq. (54). It is clear to derive it in reverse order. Ultimately we demonstrate that $\rho_{A}=\tilde{\rho}_{A}$.
@@ Appendix E The upper bound of logarithmic negativity in bPT
$\frac{1}{Z_{\sigma}}\exp\left(\sum_{k,l}(W_{\sigma})_{k,l}m_{k}m_{l}/4\right),$ (58)
@@
For particle-conserved free fermions, we can replace $\gamma$ with the Green function $G_{i,j}=\langle c_{i}c_{j}^{{\dagger}}\rangle$ and $\gamma_{\sigma}$ with $G_{\sigma}$, and demonstrate that the first two terms of Eq. (57) precisely correspond to the logarithmic negativity in uPT [38, 39, 1, 37].

### v3

@@ Appendix B Fisher-Hartwig conjecture
$\frac{\text{d}\ln\ D_{L}(\lambda)}{\text{d}\lambda}$ $=L\left[\frac{1}{\lambda-1}-\frac{k_{F}}{2\pi}\left(\frac{1}{\lambda-1}-\frac{1}{\lambda}\right)\right]-\left\{2\ln\ L+(1+\gamma_{E})+\sum_{n=1}^{\infty}\left[\frac{n}{n^{2}-\beta^{2}(\lambda)}-\frac{1}{n}\right]+\ln\ (2-2\cos k_{F})\right\}\frac{\text{d}\beta^{2}(\lambda)}{\text{d}\lambda}$ (41)
$=L\left[\frac{1}{\lambda-1}-\frac{k_{F}}{2\pi}\left(\frac{1}{\lambda-1}-\frac{1}{\lambda}\right)\right]-\frac{2\beta(\lambda)}{\pi\text{i}\lambda(\lambda-1)}\left[\ln\ L+(1+\gamma_{E})+\ln\ (2|\sin\frac{1}{2}k_{F}|)+\Upsilon(\lambda)\right],$
where $\Upsilon(\lambda)=\sum_{n=1}^{\infty}\frac{n}{n^{2}-\beta^{2}(\lambda)}-\frac{1}{n}$. For the linear term in $L$, since all of them exist as a first-order pole in $\lambda$, the integral over $\lambda$ ultimately contributes zero to them based on the residue theorem. Therefore, the entanglement quantity is given by
$\text{Entanglement quantity}=\frac{1}{\pi^{2}}\oint\text{d}\lambda f(\lambda)\frac{\beta(\lambda)}{\lambda(\lambda-1)}\left[\ln\ L+(1+\gamma_{E})+\ln\ (2|\sin\frac{1}{2}k_{F}|)+\Upsilon(\lambda)\right],$ (42)
where the integral contour is shown in Fig. 3. The residue theorem ensures that the contributions from the integration paths $\overrightarrow{FED}$ and $\overrightarrow{CAB}$ vanish. Consequently, the integral simplifies to
Entanglement quantity $=\frac{1}{\pi^{2}}\left(\int_{1+\text{i}0^{+}}^{0+\text{i}0^{+}}+\int_{0+\text{i}0^{-}}^{1+\text{i}0^{-}}\right)\text{d}\lambda f(\lambda)\frac{\beta(\lambda)}{\lambda(\lambda-1)}\left[\ln\ L+(1+\gamma_{E})+\ln\ \left(2\left|\sin\frac{1}{2}k_{F}\right|\right)+\Upsilon(\lambda)\right].$ (43)
In the complex number field the phase of $\beta\left(\lambda\right)$ must be considered to ensure a consistent and well-defined formulation
$\beta[x(\in\mathbb{R})+\mathrm{i}0^{\pm}]$ $=\frac{1}{2\pi\mathrm{i}}\left[\mathrm{Ln}\frac{\left(x-1+\mathrm{i}0^{\pm}\right)\left(x-\mathrm{i}0^{\pm}\right)}{x^{2}}\right]$ (44)
@@
$\hskip 10.00002pt+\frac{1}{\pi^{2}}\sum_{n=1}^{\infty}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}n\left\{\frac{\text{i}W(x)-\frac{1}{2}}{n^{2}-[-\text{i}W(x)+\frac{1}{2}]^{2}}+\frac{-\text{i}W(x)-\frac{1}{2}}{n^{2}-[-\frac{1}{2}-\text{i}W(x)]^{2}}+\frac{1}{n^{2}}\right\}$
@@
$E_{2}$ $=\frac{1}{\pi^{2}}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}\left(-1\right)\left[-\gamma_{E}+\sum_{n=0}^{\infty}\frac{1}{n+1}-\sum_{n=1}^{\infty}\frac{1}{n}-\frac{1}{2}\left\{\psi\left[\frac{1}{2}-\text{i}W(x)\right]+\psi\left[\frac{1}{2}+\text{i}W(x)\right]\right\}-1\right]$ (47)
$=\frac{-1}{\pi^{2}}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}\left(-\gamma_{E}-1\right)+\frac{1}{2\pi^{2}}\int_{0}^{1}\text{d}x\frac{f(x)}{x(x-1)}\left\{\psi\left[\frac{1}{2}-\text{i}W(x)\right]+\psi\left[\frac{1}{2}+\text{i}W(x)\right]\right\}.$
@@ Appendix C The eigenvalues of the density matrix after partial transpose in the overlap matrix approach
In this appendix, we will generalize our result to a random hopping model, i.e.,
$H=\sum_{i,j}t_{ij}c_{i}^{{\dagger}}c_{j}=\mathbf{c}^{{\dagger}}t\mathbf{c},$ (49)
where the elements of $t$ are random numbers following a uniform distribution while maintaining hermiticity. This can be achieved by generating random matrices $T_{1}$ and $T_{2}$ to construct $T=T_{1}+\text{i}T_{2}$, and then symmetrizing as $t=\frac{T+T^{{\dagger}}}{2}$. Other types of random $t$ can also be generated to further validate our results. Numerical findings confirm that the overlap matrix approach aligns with exact diagonalization, as shown in Fig. 7.
Figure 7: Comparison of the eigenvalues of $\rho^{T_{B}^{f}}$ and $\tilde{\rho}^{T_{B}^{f}}$: (a) the congruence of eigenvalues between $\rho^{T_{B}^{f}}$ and $\tilde{\rho}^{T_{B}^{f}}$; (b) difference in real parts of eigenvalues; and (c) difference in imaginary parts of eigenvalues. We select $N=12$ with an even bipartite geometry. The results indicate that the difference is on the order of $10^{-15}$, demonstrating near-perfect agreement [53].
Appendix D Proof of rationality of partial trace in the overlap matrix approach
In this appendix, we expand Eq. (10) into Fock space and compare with the reduced density matrix partially traced in the original basis. We prove that these two equations are completely identical.
@@ Appendix D Proof of rationality of partial trace in the overlap matrix approach
[…]
@@
$=\sum_{1\leq j_{1}<j_{2}<\cdots<j_{M}\leq N}\left[\sum_{\sigma\in S_{M}}\mathrm{sgn}(\sigma)U_{j_{\sigma(1)},1}U_{j_{\sigma(2)},2}\cdots U_{j_{\sigma(M)},M}\right]c_{j_{1}}^{\dagger}c_{j_{2}}^{\dagger}\cdots c_{j_{M}}^{\dagger}|0\rangle$
@@
where in the fourth line $\sum_{\sigma\in S_{M}}$ represents the summation over all elements of the $M$th-order symmetric group $S_{M}$ ($M$ denotes the particle number) and in the last line $J$ implies a permitted sequence of $\left\{j_{1},j_{2},\ldots,j_{M}\right\}$. Accordingly the density matrix
@@
The partial trace operation ensures that the particles in region $B$ remain the same in both bra and ket states, i.e. $J_{B}=K_{B}$. Then we decompose the numerous terms in $\rho_{A}$ based on the particles in region AA
$\rho_{A}=\sum_{m=0}^{\min(M,|A|)}\sum_{\begin{subarray}{c}J_{A},K_{A}\subseteq A\\ |J_{A}|=|K_{A}|=m\end{subarray}}\left[\sum_{\begin{subarray}{c}J_{B}\subseteq B\\ |J_{B}|=M-m\end{subarray}}\det(U_{J_{A}\cup J_{B}})\det(U_{K_{A}\cup J_{B}})^{*}\right]|J_{A}\rangle\langle K_{A}|.$ (52)
It should be noted that the maximum particle number in region $A$ is $\min\left(M,\left|A\right|\right)$, where $\left|A\right|$ represents the the number of sites in region $A$. From Sec. II, we also know that rotating $\left\{f_{\alpha}\right\}_{\alpha=1}^{M}$ to $\left\{d_{Ai}\right\}_{i=1}^{M}$ and $\left\{d_{Bi}\right\}_{i=1}^{M}$ does not alter the density matrix. Hence, the reduced density matrix can also be expressed as
$\rho_{A}=\sum_{m=0}^{\min(M,|A|)}\sum_{\begin{subarray}{c}J_{A},K_{A}\subseteq A\\ |J_{A}|=|K_{A}|=m\end{subarray}}\left\{\sum_{\begin{subarray}{c}J_{B}\subseteq B\\ |J_{B}|=M-m\end{subarray}}\det\left[(U\mathcal{U})_{J_{A}\cup J_{B}}\right]\det\left[(U\mathcal{U})_{K_{A}\cup J_{B}}\right]^{*}\right\}|J_{A}\rangle\langle K_{A}|.$ (53)
Then we expand Eq. (10):
@@
$=\prod_{\gamma=1}^{M}\left[P_{\gamma}\frac{\sum_{\alpha=1}^{M}\sum_{j\in A}c_{j}^{\dagger}U_{j,\alpha}\mathcal{U}_{\alpha,\gamma}}{\sqrt{P_{\gamma}}}|0\rangle\langle 0|\frac{\sum_{\beta=1}^{M}\sum_{j\in A}c_{j}U_{j,\beta}^{*}\mathcal{U}_{\beta,\gamma}^{*}}{\sqrt{P_{\gamma}}}+(1-P_{\gamma})|0\rangle\langle 0|\right]$
@@
where the symbol $\tilde{\rho}_{A}$ is introduced to distinguish it from the original reduced density matrix. In the fourth line we express $1-P_{\gamma}$ as $\sum_{j\in B}\sum_{\alpha,\beta=1}^{M}U_{j,\alpha}\mathcal{U}_{\alpha,\gamma}U_{j,\beta}^{*}\mathcal{U}_{\beta,\gamma}^{*}$, which can be verified by the identity $\left(\mathcal{U}^{\dagger}M_{B}\mathcal{U}\right)_{\alpha\beta}=\delta_{\alpha\beta}\left(1-P_{\alpha}\right)$. The essential identity
$0$ $=\sum_{\mu\nu}\sum_{i\in B}\mathcal{U}_{\mu\beta}^{*}U_{i\mu}^{*}U_{i\nu}\mathcal{U}_{\nu\alpha}$ (55)
@@
helps derive the equation below:
$\tilde{\rho}_{A}=\sum_{m=0}^{\min(M,|A|)}\sum_{\begin{subarray}{c}J_{A},K_{A}\subseteq A\\ |J_{A}|=|K_{A}|=m\end{subarray}}\left\{\sum_{\begin{subarray}{c}J_{B}\subseteq B\\ |J_{B}|=M-m\end{subarray}}\det\left[(U\mathcal{U})_{J_{A}\cup J_{B}}\right]\det\left[(U\mathcal{U})_{K_{A}\cup J_{B}}\right]^{*}\right\}|J_{A}\rangle\langle K_{A}|,$ (56)
from Eq. (54). It is clear to derive it in reverse order. Ultimately, we demonstrate that $\rho_{A}=\tilde{\rho}_{A}$.
@@ Appendix E The upper bound of logarithmic negativity in bPT
$\frac{1}{Z_{\sigma}}\exp\left[\sum_{k,l}\frac{(W_{\sigma})_{k,l}m_{k}m_{l}}{4}\right],$ (58)
@@
For particle-conserved free fermions, we can replace $\gamma$ with the Green’s function $G_{i,j}=\langle c_{i}c_{j}^{{\dagger}}\rangle$ and $\gamma_{\sigma}$ with $G_{\sigma}$ and demonstrate that the first two terms of Eq. (57) precisely correspond to the logarithmic negativity in uPT [37, 38, 39, 36].